# Quantum Inhomogeneous Field Theory: Unruh-Like Effects and Bubble Wall Friction

### v1

@@ 3.1 Quantum states in our background spacetime
To construct a conceptually satisfactory framework, we introduce an algebraic state, $\omega_{\rm R}^{\epsilon}$333$|0\rangle_{\rm R}^{\epsilon}$ is represented by $\omega_{\rm R}^{\epsilon}$., which is defined as a positive, normalized linear functional that assigns complex numbers to field operators. The two-point correlation function with respect to $|0\rangle_{\rm R}^{\epsilon}$ is then identified with the complex value assigned by $\omega_{\rm R}^{\epsilon}(\hat{\phi}\hat{\phi})$. Here, $\epsilon$ refers to a specific local region $x_{\epsilon}$, defined with $\epsilon$ in Figure 1. Using the Gelfand-Naimark-Segal (GNS) construction, we can return to the conventional Hilbert space representation with the algebraic state $\omega_{\rm R}^{\epsilon}$. This construction produces a representation of the algebra on a Hilbert space, $\mathcal{H}_{\omega_{\rm R}^{\epsilon}}$, where $|0\rangle_{\rm R}^{\epsilon}$ serves as the vacuum state of $\mathcal{H}_{\omega_{\rm R}^{\epsilon}}$ in our context. If the Hilbert space $\mathcal{H}_{\omega_{\rm R}^{\epsilon}}$ were to provide a consistent particle interpretation across the entire system, it would need to include states generated by scattering processes, such as transitions between massless and massive particles. However, as discussed in [26], it is strongly anticipated that, in our model governed by the quadratic Lagrangian (2.16), the Hilbert space $\mathcal{H}_{\omega_{\rm R}^{\epsilon}}$ consists of quasi-free states, which are unable to account for scattering phenomena. Furthermore, this is inconsistent with the fact that only massive particles exist in the right asymptotic region, whereas only massless particles exist in the left asymptotic region.
@@ 3.2 Hadamard regularization of the two-point function
[…]
@@
where the mode functions, $v^{(\mp)}_{k}({\bf x})$ and $(v^{(\mp)}_{k}({\boldsymbol{x}^{\prime}}))^{*}$, are given in (2.39) - (2.42) and $\omega_{k}\equiv\sqrt{k^{2}+(b\beta)^{2}}$. Here, the approximation in (3.6) represents that we have omitted the terms related with $\Phi_{\epsilon}(\boldsymbol{x})$ in (2.47). In the following, we will focus on the Wightman function of spacelike separated points ${\boldsymbol{x}}$ and ${\boldsymbol{x}}^{\prime}$, for simplicity. For timelike separated points, there is no essential difference in our result for the VEV of the energy-momentum tensor.
@@
At first, the expanded form of the Synge function444 We define $a\equiv e^{-bx_{0}}$ and perform a change of coordinates $x-x_{0}\mapsto x$. In this new coordinate system, the Ricci scalar in (2.10) is given by $\mathcal{R}=2a^{2}b^{2}e^{-bx}.$ in the background given by (2.9) is
@@
$\nabla_{\mu}\nabla_{\nu}\sigma=g_{\mu\nu}-\frac{1}{3}R_{\mu\alpha\nu\beta}\nabla^{\alpha}\sigma\nabla^{\beta}\sigma+\mathcal{O}(\sigma^{3/2}).$ (3.15)
@@ 4 Energy-momentum Tensor
$\langle T_{tt}\rangle_{\rm R}^{\epsilon}$ $=\frac{b^{2}}{16\pi}\Big{[}-\gamma+\ln 2+\frac{1}{2}\ln\Big{(}\frac{2\mu^{2}}{a^{2}b^{2}}\Big{)}\Big{]}$
$~{}~{}~{}+\frac{b^{2}}{8\pi}\Big{[}-\frac{1}{6}+\gamma-\ln 2-\frac{1}{2}\ln\Big{(}\frac{2\mu^{2}}{a^{2}b^{2}}\Big{)}\Big{]}e^{-bx}+{\cal O}\Big{(}e^{-2bx}\Big{)}\,,$ (4.6)
$\langle T_{xx}\rangle_{\rm R}^{\epsilon}$ $=\frac{b^{2}}{16\pi}\Big{[}\gamma-\ln 2-\frac{1}{2}\ln\Big{(}\frac{2\mu^{2}}{a^{2}b^{2}}\Big{)}\Big{]}$
$~{}~{}~{}+\frac{b^{2}}{8\pi}\Big{[}\frac{1}{2}-\gamma+\ln 2+\frac{1}{2}\ln\Big{(}\frac{2\mu^{2}}{a^{2}b^{2}}\Big{)}\Big{]}e^{-bx}+{\cal O}\Big{(}e^{-2bx}\Big{)}\,,$ (4.7)
@@
Since $\mathcal{R}=2a^{2}b^{2}e^{-bx}$(see footnote 4), the expansion presented above can be understood as an expansion in the weak curvature limit as $x\to\infty$.
@@
It is apparent that our expressions for the VEV of the energy-momentum tensor, as given in equations (4), (4), and (4.8), do not vanish in this limit or in Minkowski spacetime. However, this is associated with the freedom to choose the mass scale $\mu$. By selecting an appropriate mass scale $\mu$, the scheme-dependent part in the renormalization procedure can be removed. As discussed previously, the parameter $\mu$ is determined such that the VEV of the energy-momentum tensor vanishes in Minkowski spacetime.
@@
$\langle T_{tt}\rangle_{\rm M}$ $=\frac{m_{0}^{2}}{4\pi a^{2}}\left[-\gamma+\ln 2+\frac{1}{2}\ln\left(\frac{\mu^{2}}{2m_{0}^{2}}\right)\right],$
$\langle T_{xx}\rangle_{\rm M}$ $=\frac{m_{0}^{2}}{4\pi a^{2}}\left[\gamma-\ln 2-\frac{1}{2}\ln\left(\frac{\mu^{2}}{2m_{0}^{2}}\right)\right],$
@@
$\mu=\frac{m_{0}}{\sqrt{2}}e^{\gamma}.$ (4.12)
@@
$\langle T_{tt}\rangle_{\rm R}^{\epsilon}$ $=-\frac{b^{2}}{48\pi}e^{-bx}+{\cal O}\big{(}e^{-2bx}\big{)}\,,$
$\langle T_{xx}\rangle_{\rm R}^{\epsilon}$ $=\frac{b^{2}}{16\pi}e^{-bx}+{\cal O}\big{(}e^{-2bx}\big{)}\,,$ (4.13)
@@
The non-vanishing quantities in the above result represent the quantum effects of the state $|0\rangle_{{\rm R}}^{\epsilon}$ , which reflects the fact that the classical lowest energy must be zero. Note that using the result in (4.13) and recalling the expression of the Ricci scalar in footnote 4, we obtain the trace of the VEV,
@@
where $G^{+}_{\text{ren}}(\boldsymbol{x},\boldsymbol{x};\mu)=-\frac{1}{2\pi}e^{-bx}+{\cal O}\big{(}e^{-2bx}\big{)}$ as given in (3.2) with $\mu=\frac{m_{0}}{\sqrt{2}}e^{\gamma}$. At this stage, the nonvanishing trace of the VEV cannot be interpreted as a conformal anomaly but coincides with the same expression accidentally.
Comments on the $\langle T_{\mu\nu}\rangle_{{\rm R}}^{\epsilon}$-component in (4.13) are in order. Within the context of Hadamard renormalization, after eliminating the divergence from the Wightman function, the remaining term can be interpreted as the effect of a quantum state. Accordingly, using this prescription, we interpret the result $\langle T_{\mu\nu}\rangle_{\rm R}^{\epsilon}$ in (4.13) as representing the effect of the quantum state $|0\rangle_{{\rm R}}^{\epsilon}$. As discussed in section 3, the Hilbert space containing the state $|0\rangle_{{\rm R}}^{\epsilon}$ is relevant for an observer located slightly out of the right asymptotic region. It should be noted that this Hilbert space is not to be interpreted as a local one555 It is well known that there does not exist micro states for a local region. See [1, 5, 30] for example.; however, for the local observer, this Hilbert space is perceived as a “global” one. In other words, the local observer mimics the global plane wave quantization through R-quantization, as discussed in section 3, by restricting a valid range of $\omega\geq b\beta$ without imposing any constraints on position.
@@ 5.1 Quantum states in QIFT
[…]
@@ 5.3 Higgs condensate bubble expansion
$\Delta P_{\epsilon}=\frac{b^{2}}{16\pi}e^{-bx}+{\cal O}\big{(}e^{-2bx}\big{)}\,.$ (5.4)
It is clear that the positivity of the pressure difference $\Delta P_{\epsilon}$ indicates that the quantum effect of the vacuum contributes to the friction opposing the bubble wall expansion.
@@ 6 Conclusions
In this paper, we examined a free scalar field $\phi$ in a supersymmetric background metric given by (2.9) in (1+1) dimensions. After deriving the renormalized Wightman function, we explicitly obtained the covariantly conserved quantum energy-momentum tensor up to ${\cal O}\big{(}e^{-2bx}\big{)}$. Based on these results, we have shown the existence of frictional quantum effects on the expansion of the bubble wall during the electroweak phase transition in the early universe.
@@
The entire calculation in our paper follows a conceptually consistent procedure by introducing an algebraic state $\omega_{{\rm R}}^{\epsilon}$ (or $|0\rangle_{{\rm R}}^{\epsilon}$), which is locally Hadamard. Using this algebraic state, we have shown that $\langle T_{tt}\rangle_{\rm R}^{\epsilon}<0$ and $\langle T_{xx}\rangle_{\rm R}^{\epsilon}>0$. From the result $\langle T_{tt}\rangle_{\rm R}^{\epsilon}<0$, we inferred the existence of an Unruh-like effect for an observer located at $x_{\epsilon}$. Additionally, the positivity of $\langle T_{xx}\rangle_{\rm R}^{\epsilon}$ can be interpreted as a quantum frictional effect on the expansion of the bubble wall.
There are various future directions to pursue. One possibility is to extend our computations to the left asymptotic region, where a naked singularity exists. By introducing a suitable regularization parameter to handle the singularity, we can pursue similar calculations in this context. Another direction involves extending our method to a free fermion field in the framework of IFT. Additionally, exploring finite temperature effects in QIFT could provide insights relevant to the expansion of bubble walls in the early universe. Finally, investigating higher-dimensional QIFT would be of particular interest. In this case, however, the use of the Synge function in Hadamard regularization in the IFT context becomes nontrivial, as the (1+1)-dimensional conversion rule between FTCS and IFT does not apply.
@@ Acknowledgments
We appreciate conversations and discussions with Sang-A Park, as well as the valuable email correspondence with Takahiro Kubota. This work was supported by the National Research Foundation of Korea(NRF) grant with grant number RS-2023-00208047 (J.H.),  NRF-2022R1F1A1073053, RS-2023-00249608, NRF-2019R1A6A1A10073079 (O.K.), NRF-2022R1F1A1076172, NRF-2021R1A2C1003644 (S.-H.Y.), and supported by Basic Science Research Program through the NRF funded by the Ministry of Education
NRF-2020R1A6A1A03047877 (J.H. and S.-H.Y.).
@@ Appendix A Hadamard Expansion of Two-point Function
$-\frac{V_{0}}{4\pi}\nabla_{\mu}\nabla^{\mu}\ln\sigma-\frac{1}{2\pi}\left(\nabla_{\mu}V_{0}\nabla^{\mu}\sigma\right)\sigma^{-1}=\frac{1}{\sqrt{-g}}\delta(t-t^{\prime})\delta(x-x^{\prime}),$ (A.3)
$-(-\nabla_{\mu}\nabla^{\mu}+m_{0}^{2}+\xi\mathcal{R})V_{n}+2(n+1)^{2}V_{n+1}+2(n+1)\nabla_{\mu}V_{n+1}\nabla^{\mu}\sigma$
$-2(n+1)V_{n+1}\frac{1}{\sqrt{\Delta}}\nabla_{\mu}\sqrt{\Delta}\nabla^{\mu}\sigma=0,$ (A.4)
$-(-\nabla_{\mu}\nabla^{\mu}+m_{0}^{2}+\xi\mathcal{R})W_{n}+2(n+1)^{2}W_{n+1}+2(n+1)\nabla_{\mu}W_{n+1}\nabla^{\mu}\sigma$
$-2(n+1)W_{n+1}\frac{1}{\sqrt{\Delta}}\nabla_{\mu}\sqrt{\Delta}\nabla^{\mu}\sigma+4(n+1)V_{n+1}+2\nabla_{\mu}V_{n+1}\nabla^{\mu}\sigma$
$-2V_{n+1}\frac{1}{\sqrt{\Delta}}\nabla_{\mu}\sqrt{\Delta}\nabla^{\mu}\sigma=0$ (A.5)

### v4

@@ 3.1 Quantum states in our background spacetime
[…]
@@ 3.2 Hadamard regularization of the two-point function
$G^{+}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\int d\mu_{\eta}\,u_{\eta}({\boldsymbol{x}})u^{*}_{\eta}({\boldsymbol{x}}^{\prime})\,$, (3.3)
@@
where the mode functions, $v^{(\mp)}_{k}({\bf x})$ and $(v^{(\mp)}_{k}({\boldsymbol{x}^{\prime}}))^{*}$, are given in (2.44) - (2.47) and $\omega_{k}\equiv\sqrt{k^{2}+(b\beta)^{2}}$. Here, the approximation in (3.6) represents that we have omitted the terms related with $\Phi_{\epsilon}(\boldsymbol{x})$ in (2.52). In the following, we will focus on the Wightman function of spacelike separated points $\boldsymbol{x}$} and $\boldsymbol{x}$$^{\prime}$, for simplicity. Although we use a timelike point-splitting, the VEV of the energy-momentum tensor remains unchanged after applying the Hadamard renormalization prescription for a spacelike point-splitting.
@@
At first, the expanded form of the Synge function [^6] See the footnote 2 for the parameter $a$. in the background given by (2.9) is
@@
$\nabla_{\mu}\nabla_{\nu}\sigma=g_{\mu\nu}-\frac{1}{3}\mathcal{R}_{\mu\alpha\nu\beta}\nabla^{\alpha}\sigma\nabla^{\beta}\sigma+\mathcal{O}(\sigma^{3/2}).$ (3.15)
@@ 4 Energy-momentum Tensor
$\langle T_{tt}\rangle_{\rm R}^{\epsilon}$ $=\frac{b^{2}}{16\pi}\Big[\frac{1}{2}-\gamma+\ln 2+\frac{1}{2}\ln\Big(\frac{2\mu^{2}}{a^{2}b^{2}}\Big)\Big]\Big[1-2e^{-bx}\Big]$
$~~~-\frac{b^{2}}{12\pi}e^{-bx}+{\cal O}\Big(e^{-2bx}\Big)\,$, (4.6)
$\langle T_{xx}\rangle_{\rm R}^{\epsilon}$ $=\frac{b^{2}}{16\pi}\Big[-\frac{1}{2}+\gamma-\ln 2-\frac{1}{2}\ln\Big(\frac{2\mu^{2}}{a^{2}b^{2}}\Big)\Big]\Big[1-2e^{-bx}\Big]+{\cal O}\Big(e^{-2bx}\Big)\,$, (4.7)
@@
Since $\mathcal{R}=2a^{2}b^{2}e^{-bx}$(see footnote 2), the expansion presented above can be understood as an expansion in the weak curvature limit as $x\to\infty$.
@@
It is apparent that our expressions for the VEV of the energy-momentum tensor, as given in equations (4), (4.7), and (4.8), do not vanish in this limit or in Minkowski spacetime. However, this is associated with the freedom to choose the mass scale $\mu$. By selecting an appropriate mass scale $\mu$, the scheme-dependent part in the renormalization procedure can be removed. As discussed previously, the parameter $\mu$ is determined such that the VEV of the energy-momentum tensor vanishes in Minkowski spacetime.
@@
$\langle T_{tt}\rangle_{\rm M}$ $=\frac{m_{0}^{2}}{4\pi a^{2}}\left[\frac{1}{2}-\gamma+\ln 2+\frac{1}{2}\ln\!\left(\frac{\mu^{2}}{2m_{0}^{2}}\right)\right],$
$\langle T_{xx}\rangle_{\rm M}$ $=\frac{m_{0}^{2}}{4\pi a^{2}}\left[-\frac{1}{2}+\gamma-\ln 2-\frac{1}{2}\ln\!\left(\frac{\mu^{2}}{2m_{0}^{2}}\right)\right],$
@@
$\mu=\frac{m_{0}}{\sqrt{2}}\,e^{\gamma-\frac{1}{2}}.$ (4.12)
@@
$\langle T_{tt}\rangle_{\rm R}^{\epsilon}$ $=-\frac{b^{2}}{12\pi}e^{-bx}+{\cal O}\big(e^{-2bx}\big)\,$,
$\langle T_{xx}\rangle_{\rm R}^{\epsilon}$ $={\cal O}\big(e^{-2bx}\big)\,$, (4.13)
@@
The non-vanishing quantities in the above result represent the quantum effects of the state $|0\rangle_{{\rm R}}^{\epsilon}$ , which reflects the fact that the classical lowest energy must be zero. Note that using the result in (4.13) and recalling the expression of the Ricci scalar in footnote 2, we obtain the trace of the VEV,
@@
where $G^{+}_{\text{ren}}(\boldsymbol{x},\boldsymbol{x};\mu)=-\frac{1}{2\pi}e^{-bx}+{\cal O}\big(e^{-2bx}\big)$ as given in (3.2) with $\mu=\frac{m_{0}}{\sqrt{2}}\,e^{\gamma-\frac{1}{2}}$. At this stage, the nonvanishing trace of the VEV cannot be interpreted as a conformal anomaly but coincides with the same expression accidentally.
Comments on the $\langle T_{\mu\nu}\rangle_{{\rm R}}^{\epsilon}$-component in (4.13) are in order. Within the context of Hadamard renormalization, after eliminating the divergence from the Wightman function, the remaining term can be interpreted as the effect of a quantum state. Accordingly, using this prescription, we interpret the result $\langle T_{\mu\nu}\rangle_{\rm R}^{\epsilon}$ in (4.13) as representing the effect of the quantum state $|0\rangle_{{\rm R}}^{\epsilon}$. As discussed in section 3, the Hilbert space containing the state $|0\rangle_{{\rm R}}^{\epsilon}$ is relevant for an observer located slightly out of the right asymptotic region. It should be noted that this Hilbert space is not to be interpreted as a local one [^7] It is well known that there does not exist micro states for a local region. See [1, 5, 30] for example.; however, for the local observer, this Hilbert space is perceived as a “global” one. In other words, the local observer mimics the global plane wave quantization through R-quantization, as discussed in section 3, by restricting a valid range of $\omega\geq b\beta$ without imposing any constraints on position.
@@ 5.1 Quantum states in QIFT
As shown in subsection 2.3, while the L-vacuum $|0\rangle_{\rm L}$ defined in (2.43) is suitable for the left region $x\to$-$\infty$, it is not an appropriate candidate for the vacuum of the scalar field $\phi$ in the right region $x\to\infty$. Since (massless) particles of the energy $0\leq\omega\leq b\beta$ in the left region do not exist in the right asymptotic region, the L-vacuum cannot be an appropriate vacuum for the quantum field in the right region, $x\to+\infty$. For similar reasons, the R-vacuum $|0\rangle_{\rm R}$ is not a good candidate for a vacuum state in the left asymptotic region. As discussed in [26], within the L-quantization scheme, the Fock space $\mathcal{F}_{\mathrm{L}}$ is constructed by massless particles, whereas the R-quantization scheme yields a Fock space $\mathcal{F}_{\mathrm{R}}$ consisting of massive particles. The Fock space $\mathcal{F}_{\mathrm{L}}$ provides a good representation in the region where $x\to$-$\infty$, while $\mathcal{F}_{\mathrm{R}}$ offers a valid description for $x\to\infty$. These two Fock spaces cannot be connected by a unitary transformation, making the existence of a global Hadamard state implausible.
@@ 5.3 Higgs condensate bubble expansion
$\Delta P_{\epsilon}={\cal O}\big(e^{-2bx}\big)\,.$ (5.4)
From the point of view the inhomogeneous field theory, the quantum vacuum yields an Unruh–like contribution as the dominant term, scaling as $\mathrm{e}^{-bx}$. However, any pressure—whether tending to hinder or to assist the bubble–wall expansion—does not arise at this leading order. Accordingly, the influence of the pressure difference $\Delta P$ enters only at the next–to–leading order, suppressed by at least $\mathrm{e}^{-2bx}$.
@@ 6 Conclusions
In this paper, we examined a free scalar field $\phi$ in a supersymmetric background metric given by (2.9) in (1+1) dimensions. After deriving the renormalized Wightman function, we explicitly obtained the covariantly conserved quantum energy-momentum tensor up to ${\cal O}\big(e^{-2bx}\big)$. Based on these results, we have shown the existence of the Unrhu-like quantum effects in the leading order $(e^{-bx})$on the expansion of the bubble wall during the electroweak phase transition in the early universe.
@@
The entire calculation in our paper follows a conceptually consistent procedure by introducing an algebraic state $\omega_{{\rm R}}^{\epsilon}$ (or $|0\rangle_{{\rm R}}^{\epsilon}$), which is locally Hadamard. Using this algebraic state, we have shown that $\langle T_{tt}\rangle_{\rm R}^{\epsilon}<0$ and $\langle T_{xx}\rangle_{\rm R}^{\epsilon}=0$ in the leading order $e^{-bx}$. From the result $\langle T_{tt}\rangle_{\rm R}^{\epsilon}<0$, we inferred the existence of an Unruh-like effect for an observer located at $x_{\epsilon}$.
There are various future directions to pursue. One possibility is to extend our computations to the left asymptotic region, where a naked singularity exists. By introducing a suitable regularization parameter to handle the singularity, we can pursue similar calculations in this context. Another direction involves extending our method to a free fermion field in the framework of IFT [46]. Additionally, exploring finite temperature effects in QIFT could provide insights relevant to the expansion of bubble walls in the early universe. Finally, investigating higher-dimensional QIFT would be of particular interest. In this case, however, the use of the Synge function in Hadamard regularization in the IFT context becomes nontrivial, as the (1+1)-dimensional conversion rule between FTCS and IFT does not apply.
@@ Acknowledgments
We appreciate conversations and discussions with Sang-A Park, as well as the valuable email correspondence with Takahiro Kubota. This work was supported by the National Research Foundation of Korea (NRF) grant with grant number RS-2023-00208047 (J.H.), RS-2019-NR040081, RS-2023-00249608 (O.K.), NRF-2022R1F1A1076172, NRF-2021R1A2C1003644 (S.-H.Y.), and supported by Basic Science Research Program through the NRF funded by the Ministry of Education NRF-2020R1A6A1A03047877 (J.H. and S.-H.Y.).
@@ Appendix A Hadamard Expansion of Two-point Function
$\frac{V_{0}}{4\pi}\Box\ln\sigma+\frac{1}{2\pi}\left(\nabla_{\mu}V_{0}\nabla^{\mu}\sigma\right)\sigma^{-1}=-\frac{1}{\sqrt{-g}}\delta(t-t^{\prime})\delta(x-x^{\prime}),$ (A.3)
$(-\Box+m_{0}^{2}+\xi\mathcal{R})V_{n}-2(n+1)^{2}V_{n+1}-2(n+1)\nabla_{\mu}V_{n+1}\nabla^{\mu}\sigma$
$~~+2(n+1)V_{n+1}\frac{1}{\sqrt{\Delta}}\nabla_{\mu}\sqrt{\Delta}\nabla^{\mu}\sigma=0,$ (A.4)
$(-\Box+m_{0}^{2}+\xi\mathcal{R})W_{n}-2(n+1)^{2}W_{n+1}-2(n+1)\nabla_{\mu}W_{n+1}\nabla^{\mu}\sigma$
$~~+2(n+1)W_{n+1}\frac{1}{\sqrt{\Delta}}\nabla_{\mu}\sqrt{\Delta}\nabla^{\mu}\sigma-4(n+1)V_{n+1}-2\nabla_{\mu}V_{n+1}\nabla^{\mu}\sigma$
$~~+2V_{n+1}\frac{1}{\sqrt{\Delta}}\nabla_{\mu}\sqrt{\Delta}\nabla^{\mu}\sigma=0$ (A.5)
@@
Appendix B Order Estimate of $\Phi_{\epsilon}$
In this Appendix, we present an argument to estimate the order of $\Phi_{\epsilon}$ in (2.53). To proceed, let us consider a “position-dependent” dispersion relation given by
$\omega^{2}(x)=k^{2}+m^{2}(x)\,.$ (B.1)
Under the condition $\beta=\frac{1}{2}=2\xi$, we expand the mass function around the point $x=x_{\epsilon}$, which is located near $x=\infty$. This expansion yields
$m^{2}(x)=\left(\frac{m_{0}}{a}\right)^{2}+e^{-bx}\left[\frac{b^{2}}{2a}-2\frac{m_{0}^{2}}{a^{3}}\right]+e^{-2bx}\left[-\left(\frac{b}{a}\right)^{2}+3\frac{m_{0}^{2}}{a^{4}}\right]+{\cal O}\left(e^{-3bx}\right)\,.$ (B.2)
In our case, the relation $ab=2m_{0}$ holds, which leads to
$\frac{b^{2}}{2a}-2\frac{m_{0}^{2}}{a^{3}}=0\,.$
Consequently, the expansion of the mass function simplifies to
$m^{2}(x)=(b\beta)^{2}-\frac{b^{2}}{4a^{2}}\,e^{-2bx}+{\cal O}\left(e^{-3bx}\right)\,.$ (B.3)
Now, using the dispersion relation, we deduce the relation between $x_{\epsilon}$ and $\epsilon$. As illustrated in Fig. 1, the energy at the position $x=x_{\epsilon}$ for the lowest momentum (i.e., $k=0$) is given by
$\omega=b\beta-\epsilon\qquad(\epsilon\ll b\beta)\,.$
Substituting this into the dispersion relation (B.1), we obtain
$\omega^{2}=(b\beta-\epsilon)^{2}=m^{2}(x_{\epsilon})=\left(\frac{b}{2}\right)^{2}-\frac{b^{2}}{4a^{2}}\,e^{-2bx_{\epsilon}}+{\cal O}\left(e^{-3bx_{\epsilon}}\right)\,.$ (B.4)
Expanding the left-hand side for small $\epsilon$, and recalling $\beta=\frac{1}{2}$, we find
$\frac{1}{2b\beta}\epsilon=\frac{1}{b}\epsilon=\frac{b}{4a^{2}}e^{-2bx_{\epsilon}}+{\cal O}\left(e^{-3bx_{\epsilon}}\right)\,.$ (B.5)
Recalling $W\sim e^{-i(\omega t\pm kx)}\sim{\cal O}(1)$ in (2.53), we can estimate the order of the integral term in $\Phi_{\epsilon}$ as
$\int^{b\beta}_{b\beta-\epsilon}d\omega\left(\cdots\right)\simeq\epsilon\times{\cal O}(1)\simeq{\cal O}\left(e^{-2bx_{\epsilon}}\right)\,.$ (B.6)
Since this term is of higher order compared to the terms of interest, which are up to ${\cal O}\left(e^{-bx_{\epsilon}}\right)$ in $\Phi_{\epsilon}$, we conclude that the integral term over the range $b\beta-\epsilon<\omega<b\beta$ in $\Phi_{\epsilon}$ can be neglected in our analysis. Consequently, the contribution of $\Phi_{\epsilon}$ to the 2-point function is ${\cal O}(e^{-2bx_{\epsilon}})$, which is negligible to the leading order of interest in the 2-point function, ${\cal O}(e^{-bx_{\epsilon}})$.